EVA Report – February 23rd

Crew 222 EVA Report 23-02-2020
EVA # 7
Author: Nicolas Bourdeaud (HSO)
Purpose of EVA: Replacement of the battery of MegaAres.
Start time: 4:53pm
End time: 5:26pm

Narrative: Everything went just fine. We rapidly changed the battery and
assure the waterproofness of the experiment, despite the mud all around
the experiment.

Destination: Experiments Site.
Coordinates (use UTM NAD27 CONUS): 12S 0518494-4250679
Participants: Nicolas Bourdeaud (HSO, EVA Leader), Nina Sedbon (Scientist)
Road(s) and routes per MDRS Map: Entrance road – Cow Dung Road
Mode of travel: Rover – Curiosity

Astronomy Report – February 23rd

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/23/2020

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: As weather has improved, the active
observations should progress tonight.

Images submitted with this report: Taking advantage of bad weather, the
M101 was processed again, taking Peter’s tips into account. The bad
pixels are less visible now. Sorry I had only saved in PNG, not in JPEG!

Problems Encountered: None

MUSK OBSERVATORY

Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Operations Report – February 23rd

Crew 222 Operations Report 02/23/2020

SOL: 7

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Friday generator:
Off at 5:30am
On at 5:00pm SOC was 80%

Saturday generator:
Off at 6:30am
On at 9:00pm SOC was XX%

Diesel Reading – 74%

Station Propane Reading – 30% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149426.79 units

Water (static tank) – 319 gallons

Static to Loft Pump used – yes

Water in GreenHab – 80 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours: 146.2
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.4
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.4
Beginning charge: 100%
Ending charge: 90%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s
parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All was fixed yesterday evening. All ok now.

Summary of suits and radios: Suits 8 and 9 used and radio tested for the
EVA. all ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Journalist Report – February 23rd

Crew 222 Journalist Report 23Feb2020

Author: Marie Bochard, Crew Journalist

Sol 7

Title: White land

It’s snowing outside! Here, the temperature easily drops below 0°C, staying at about -12°C for the majority of the night. The landscape feels changed, and it’s lucky that this change happened today since it is the middle of our mission. We started the simulation on Sunday, it’s been exactly one week since we started the mission, and there is one week left. We now have new sights to discover and new areas to explore in a different light!

Today is rest day. We’ve been working hard and going on EVA every day for a week, which has been tiring. We focused on ourselves today. We had new discussions, we read new books, we started putting together a summary video of our trip to get here… and of course caught up on sleep. Waking up at 6:45 to do sports every morning isn’t easy, even if it is worth it.

At the end of the mission, I can assure you that everyone will be able to solve a Rubik’s cube, thanks to Benjamin’s help! We also discovered space chess and battled hard.

We also took time to reflect on our mission so far. The beauty of the landscape is unanimously what we will remember most. Living with and only seeing six other people for a week has been a new experience for the majority of us, and so far, so good! We get along great.

Cooking has brought us closer: we bake bread daily, and sharing recipes between crew members makes us discover new tastes… and cooking abilities.

GreenHab Report – February 23rd

Crew 222 GreenHab Report – 23/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Ambient with door open

Average temperatures: 23°C

Hours of supplemental light: NA

Daily water usage for crops: 9 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 80 gallons

Time(s) of watering crops: 2

Changes to crops: NA

Narrative:Today, after watering plants, I exposed watercress in gelatin
to UV or/and notes of music again. I did watered plants again
tonight but nothing special happened in the rest of the day.

Harvest: NA

Support/supplies needed: As I expected, today I had to use the blue tap
so i don’t know know if i’ll have enaf water in the blue tank to
water plants until the end of the simulation

Sol Summary – February 23rd

Crew 222 Sol Summary Report 23-02-2020
Sol:7
Summary Title: Calm day in the Hab
Author’s name: Benjamin Auzou
Mission Status: All things running
Sol Activity Summary:
Today was a calm day in the Hab, with a small EVA in the end of the
afternoon. We spend a lot of time talking and debating about various topics.
This was also the last session of one of our human factors experiments.

Look Ahead Plan:
Tomorrow, we will get back to the habitual planning, and the experiments

Anomalies in work: none
Weather: very muddy
Crew Physical Status: the crew member suffering from the shoulder is not
better nor worse, the rest of the crew is safe and sound.
EVA: MegaARES battery replacement
Reports to be filed: Journalist report, Operations Report, GreenHab
Report, Astronomy Report, EVA Report, Photo report, Photo of the Day
Support Requested: Trashes in the engineering airlock